OpenStudy (anonymous):

Simplify 8.2(2.1 x 106) and write in scientific notation

OpenStudy (inkyvoyd):

man could you have more conversations about bullpellet nothing, instead of asking for math help?

OpenStudy (bibby):

\(a*b*c=a*c*b\) use this property and rewrite it as \(8.2(2.1 \times 10^6)=(8.2)(2.1)(10^6)=(8.2*2.1)(10^6)\)

OpenStudy (anonymous):

that literally didnt help witth anything, thanks though.

OpenStudy (bibby):

wow you don't have to be rude, dude what is 8.2*2.1?

OpenStudy (anonymous):

Dude , i said thanks ?! there was nothin rude in that reply !

OpenStudy (bibby):

you're not even trying. what is 8.2*2.1? "that literally didnt help witth anything"

OpenStudy (anonymous):

is it 10^6 or 106

OpenStudy (anonymous):

10^6

OpenStudy (anonymous):

8.2(2.1×10^6) well it become 17.22 x 10^6 which can be writtern as 1.722 x10^7

OpenStudy (bibby):

\(8.2(2.1 \times 10^6) =(8.2 \times 2.1)10^6=17.22\times10^6\)